unseen:

Ted weighs twice as much as Julie. Mike weighs three times as much as Julie. Together, Ted, Mike, and Julie weigh 210 lbs. What is the weight of each person?

iosangel:

K. so, first off lets make julies weight be x and teds weight be 2x and mikes weight be 3x and all of them together weigh 210. x+2x+3x=210. now lets solve. so firsts off we combine like terms like this: (x+2x+3x)=210= 6x=210 Now we divide both sides by 6 (and the 6x will be crossed out) and 210/6 is 35 now we know that julies weight is 35 and teds weight is 2x so 2*35=70 so teds weight is 70 now mikes weight: so 3*35=105<-- mikes weight

iosangel:

Oh and i forgot one more thing. CHECK IT! so if we add all their weights itll add up to 210 105+70+35=210

iosangel:

opss i meant to type (and the 6x/6 will be crossed out)

unseen:

its right @iosangel

iosangel:

e.e Great to know!